OpenStudy (anonymous):

WILL MEDALL & FANN Solve the equation. 1/2x+14 - 9/x+7=-6 A. x=-101/12 B. x= 101/12 C. x= -67/12 D. x= 67/12

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HEY!!

OpenStudy (anonymous):

Heeyy

OpenStudy (misty1212):

\[\frac{1}{2x+14}-\frac{9}{x+7}=-6\]?

OpenStudy (anonymous):

how do i solve it?

OpenStudy (misty1212):

the least common multiple of the denominators is \(2x+14\)

OpenStudy (misty1212):

multiply both sides by \(2x+14\) to cancel the denominators and get \[1-18=-6(2x+14)\]

OpenStudy (misty1212):

then solve like any other linear equation \[-17=-12x-84\] etc etc

OpenStudy (misty1212):

you good from there?

OpenStudy (anonymous):

just solve that ^^ and thats it?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

okay i got it :) thxxx !!! :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]